# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ 3.1 Model
We consider a nonparametric binary regression problem in which logit-transformed probabilities of binary responses have a logistic-beta process prior. We highlight advantages of this prior in simplifying posterior inference and prediction, while facilitating extensions. Let $z(x)\in\{0,1\}$ be binary data indexed by $x\in\mathscr{X}$. To induce a $\mathrm{Beta}(a,b)$ prior for $\operatorname*{pr}\{z(x)=1\}$ for each $x$, while enabling inference on probabilities at arbitrary $x^{*}$ values, we choose
$z(x_{i})\mid\eta(x_{i})$ $\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]\quad(i=1,\dots,n),\quad\eta(\cdot)\sim{\textsc{lbp}}(a,b,\mathcal{R}).$ (8)
By logistic-beta properties, the prior probabilities are marginally $\mathrm{Beta}(a,b)$ distributed, providing a natural extension from the independent beta-Bernoulli model to a dependent process model. The right panel of Fig. 3 shows an example of posterior prediction. We simulated 600 binary data points with $\operatorname*{pr}\{z(x)=1\}=\sigma\{\cos(\pi x)\}$, and fitted the latent logistic-beta process model (8) with $a=2,b=4$ and Matérn correlation kernel with fixed parameters $(\varrho,\nu)=(0.3,1.5)$. The middle part is slightly biased towards zero and has a wider credible interval, reflecting shrinkage towards the prior and uncertainty due to sparse data.
Figure 3: Illustration of the logistic-beta process on $\mathscr{X}=[0,3]$ with Matérn correlation kernel. (Left) Eight realizations of $\sigma\{\eta(x)\}$ with $\eta(\cdot)\sim{\textsc{lbp}}(2,4,\mathcal{R}_{M})$, following $\mathrm{Beta}(2,4)$ marginally. (Right) Posterior mean (solid) and 95% credible interval of latent success probabilities given the binary data $z(x_{i})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]$, $i=1,\dots,n$ represented as small vertical bars. The dashed line corresponds to the true success probabilities.
@@ 3.2 Posterior computation
We describe a Markov chain Monte Carlo algorithm for model (8). The key components are Pólya-Gamma data augmentation (Polson et al.,, 2013) and the normal variance-mean mixture representation (5), which leads to a latent Gaussian model conditional on the Pólya mixing variable $\lambda$. In Algorithm 9, we describe a blocked Gibbs sampler. We use notations $z_{1:n}=\{z(x_{1}),\dots,z(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$, $\omega=(\omega_{1},\dots,\omega_{n})^{\mathrm{\scriptscriptstyle T}}$ and $\Omega=\hbox{diag}(\omega)$ for Pólya-Gamma variables, $\eta=(\eta_{1},\dots,\eta_{n})^{\mathrm{\scriptscriptstyle T}}=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ and $R=\{\mathcal{R}(x_{i},x_{j})\}_{i,j}$. Here we assume that the correlation kernel $\mathcal{R}$ is known for conciseness, but incorporating unknown correlation kernel parameters is straightforward via an additional sampling step.
[1] Sample $\omega_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\pi(\omega_{i}\mid\eta_{i})$ for $i=1,\dots,n$, which is Pólya-Gamma with parameters $1$ and $\eta_{i}$.
[2] Sample $\lambda$ from $\pi(\lambda\mid\omega,z_{1:n})$, where $\eta$ is marginalized out,
$\pi(\lambda\mid\omega,z_{1:n})\propto{\pi_{\textsc{po}}}(\lambda;a,b)\times N_{n}\big{\{}\Omega^{-1}(z_{1:n}-0.51_{n});0.5\lambda(a-b)1_{n},\lambda R+\Omega^{-1}\big{\}}.$ (9)
[3] Sample $\eta$ from $\pi(\eta\mid\lambda,\omega,z_{1:n})\propto\pi(z_{1:n}\mid\eta,\omega)\times\pi(\eta\mid\lambda)$, which is multivariate normal.
@@
Step 1 of Algorithm 9 updates Pólya-Gamma auxiliary variables $\omega$ using a rejection sampler (Polson et al.,, 2013). Data augmentation with Pólya-Gamma auxiliary variables converts the Bernoulli likelihood (8) into a normal likelihood in terms of logit probabilities $\eta$ conditional on Pólya-Gamma variables $\omega$. Leveraging (5), $\pi(\eta\mid\lambda)$ is also a multivariate normal in terms of $\eta$. This leads to joint conditional $\pi(\eta,\lambda\mid\omega,z_{1:n})\propto\pi(z_{1:n}\mid\eta,\omega)\pi(\eta\mid\lambda)\pi(\lambda)$ that is multivariate normal in terms of $\eta$. Thus, the computation of collapsed conditional $\pi(\lambda\mid\omega,z_{1:n})=\int\pi(\lambda,\eta\mid\omega,z_{1:n})\mathrm{d}\eta$ can integrate out $\eta$ by exploiting normal-normal conjugacy, greatly improving mixing behavior. Furthermore, again by conjugacy, Step 3 of Algorithm 9 updates $\eta$ from a multivariate normal; for example when correlation matrix $R$ is full rank,
$\eta\mid\omega,\lambda,z_{1:n}\sim N_{n}\big{[}(\Omega+\lambda^{-1}R^{-1})^{-1}\{(z_{1:n}-0.51_{n})+0.5(a-b)R^{-1}1_{n}\},(\Omega+\lambda^{-1}R^{-1})^{-1}\big{]}.$
Thus, the logistic-beta process offers significant computational benefits compared to multivariate beta constructions that rely on coordinate-wise updating (Trippa et al.,, 2011; DeYoreo and Kottas,, 2018) or Metropolis-Hastings in high-dimensional space (Arbel et al.,, 2016). Sampling $\lambda$ from a collapsed conditional density in Step 2 remains nontrivial and involves the Pólya density ${\pi_{\textsc{po}}}$ that is numerically unstable to evaluate; in Appendix A we propose a novel class of adaptive Pólya proposals that bypass evaluation of the Pólya density based on Proposition 1. In Appendix C, we present further computational strategies based on scalable Gaussian process methods that preserve marginal variances, including normalized feature map kernel with $q\ll n$; the representations (6) and (7) both highlight the dimension reduction from $n$ to $q$.
@@ 4 Logistic-beta dependent Dirichlet processes
4.1 Definition and prior properties
@@
The latent logistic-beta process model can be easily employed in more complex Bayesian nonparametric models. For example, we introduce a logistic-beta dependent Dirichlet process for modeling dependent random probability measures. Random probability measures $\{G_{x}:x\in\mathscr{X}\}$ follow a logistic-beta dependent Dirichlet process with concentration $b$, correlation kernel $\mathcal{R}$, and independent atom processes $\{\theta_{h}(x):x\in\mathscr{X}\}$, $h\in\mathds{N}$ with $G^{0}_{x}$ the base measure, if
@@
$G_{x}(\cdot)=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x)\}\prod_{l<h}\left[1-\sigma\{\eta_{l}(x)\}\right]\Big{)}\delta_{\theta_{h}(x)}(\cdot),\,\,\,\,\eta_{h}(\cdot)\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R})\,\,(h\in\mathds{N}).$ (10)
@@
Since the stick-breaking ratios $\sigma\{\eta_{h}(x)\}$ have $\mathrm{Beta}(1,b)$ marginals for any $x\in\mathscr{X}$ and are independent across $h=1,2\dots$, the random probability measure $G_{x}$ for any given $x\in\mathscr{X}$ marginally follows a Dirichlet process with concentration parameter $b$. The hyperparameter $b$ has a familiar interpretation, aiding in elicitation.
To gain a deeper understanding of the dependence structure induced by the logistic-beta dependent Dirichlet process, we study the correlation between random probability measures, focusing on the dependence induced by the weights by assuming an atom process that does not depend on $x$, referred to as a single-atoms process.
Theorem 1.
@@
Consider a single-atoms logistic-beta dependent Dirichlet process $\{G_{x}:x\in\mathscr{X}\}$ with concentration parameter $b$ and correlation kernel $\mathcal{R}$, where its atoms are identically and independently distributed from a nonatomic base measure. Let $\mu(x_{i},x_{j})=E[\sigma\{\eta(x_{i})\}\sigma\{\eta(x_{j})\}]$ with $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$. Then, for any Borel set $B$,
$\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}={(1+b)^{2}}/\left\{2\mu(x_{i},x_{j})^{-1}-(1+b)\right\}.$ (11)
The correlation (11) does not depend on the choice of $B$ and attains a maximum of $1$ when $x_{i}=x_{j}$. The minimum correlation is achieved when $\mu(x_{i},x_{j})$ is minimized, whose value depends on $b$ due to the nontrivial correlation lower bound described in Corollary 1.
To investigate the range of dependence, we analyze and compare the greatest lower bounds of correlation (11) under four different single-atoms dependent Dirichlet process models based on dependent beta stick-breaking ratios: (M1) logistic-beta dependent Dirichlet process; (M2) The process proposed by DeYoreo and Kottas, (2018); (M3) The process proposed by Nieto-Barajas et al., (2012); (M4) Copula-based dependent Dirichlet process. For M4, the minimum correlation is achieved with counter-monotonic beta random variables; examples include Gaussian copula with perfect negative correlation. Under the single-atoms processes, where atoms are shared across $x$, the correlation (11) cannot be negative even though the correlation between weights is negative (Ascolani et al.,, 2023). For completeness, we add brief descriptions of M2–M4 and their correlation lower bounds in Appendix B.3.
@@
Figure 4: Range of dependence induced by single-atoms dependent Dirichlet processes (M1)–(M4) against the concentration parameter $b$ on a $\log_{2}$ scale. (Left) Infimum of $\operatorname*{corr}\{V_{h}(x_{i}),V_{h}(x_{j})\}$, the correlation between $\mathrm{Beta}(1,b)$ stick-breaking variables. (Right) Infimum of $\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$, the correlation between random probability measures evaluated on the same set. All processes (M1)–(M4) attain perfect correlation of 1 when $x_{i}=x_{j}$.
Figure 4 illustrates the range of dependence that processes M1–M4 can capture, with M4 being the widest. The behaviour of the minimum correlation of M1 varies distinctly under different values of $b$ compared to others. For small to moderate $b$, M1 can capture a much wider range of dependence compared to M2 and M3 by allowing negatively correlated stick-breaking ratios. When $b$ is large, the correlation range of M1 becomes limited due to the nontrivial lower bound of correlation described in Corollary 1, and it gives useful guidance on the choice of $b$. In practice, we suggest choosing $b$ less than 8 for M1 so that a negative correlation between beta stick-breaking variables is possible.

### v3

@@ 3.1 Model
We consider a nonparametric binary regression problem in which logit-transformed probabilities of binary responses have a logistic-beta process prior. We highlight advantages of this prior in simplifying posterior inference and prediction while facilitating extensions. Let $z(x)\in\{0,1\}$ be binary data indexed by $x\in\mathscr{X}$. To induce a $\mathrm{Beta}(a,b)$ prior for $\operatorname*{pr}\{z(x)=1\}$ for each $x$, while enabling inference on probabilities at arbitrary $x^{*}$ values, we choose
$z(x_{i})\mid\eta(x_{i})$ $\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]\quad(i=1,\dots,n),\quad\eta(x)\sim{\textsc{lbp}}(a,b,\mathcal{R}).$ (8)
By logistic-beta properties, the prior probabilities are marginally $\mathrm{Beta}(a,b)$ distributed, providing a natural extension from the independent beta-Bernoulli model to a dependent process model. The right panel of Figure 3 shows an example of posterior prediction. We simulated 600 binary data points with $\operatorname*{pr}\{z(x)=1\}=\sigma\{\cos(\pi x)\}$, and fitted the latent logistic-beta process model (8) with $a=2,b=4$ and Matérn correlation kernel with fixed parameters $(\varrho,\nu)=(0.3,1.5)$. The middle part is slightly biased towards zero and has a wider credible interval, reflecting shrinkage towards the prior and uncertainty due to sparse data.
@@ 3.2 Posterior computation
We describe a Markov chain Monte Carlo algorithm for model (8). The key components are Pólya-Gamma data augmentation (Polson et al.,, 2013) and the normal variance-mean mixture representation (5), which leads to a latent Gaussian model conditional on the Pólya mixing variable $\lambda$. In Algorithm 9, we describe a blocked Gibbs sampler. We use notations $\mathbf{z}=\{z(x_{1}),\dots,z(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$, $\bm{\omega}=(\omega_{1},\dots,\omega_{n})^{\mathrm{\scriptscriptstyle T}}$ and $\bm{\Omega}=\hbox{diag}(\bm{\omega})$ for Pólya-Gamma variables, $\bm{\eta}=(\eta_{1},\dots,\eta_{n})^{\mathrm{\scriptscriptstyle T}}=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ and $\mathbf{R}=\{\mathcal{R}(x_{i},x_{j})\}_{i,j}$. Here we assume that the correlation kernel $\mathcal{R}$ is known for conciseness, but incorporating unknown correlation kernel parameters is straightforward via an additional sampling step.
[1] Sample $\omega_{i}\mid\eta_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{PG}(1,\eta_{i})$ for $i=1,\dots,n$, the Pólya-Gamma distribution.
[2] Sample $\lambda$ from $\pi(\lambda\mid\bm{\omega},\mathbf{z})$, where $\bm{\eta}$ is marginalized out,
$\pi(\lambda\mid\bm{\omega},\mathbf{z})\propto{\pi_{\textsc{po}}}(\lambda;a,b)\times N_{n}\big{\{}\bm{\Omega}^{-1}(\mathbf{z}-0.5\bm{1}_{n});0.5\lambda(a-b)\bm{1}_{n},\lambda\mathbf{R}+\bm{\Omega}^{-1}\big{\}}.$ (9)
[3] Sample $\bm{\eta}$ from $\pi(\bm{\eta}\mid\lambda,\bm{\omega},\mathbf{z})\propto\pi(\mathbf{z}\mid\bm{\eta},\bm{\omega})\times\pi(\bm{\eta}\mid\lambda)$, which is multivariate normal.
@@
Step 1 of Algorithm 9 updates Pólya-Gamma auxiliary variables $\bm{\omega}$ using a rejection sampler (Polson et al.,, 2013). Data augmentation with Pólya-Gamma auxiliary variables converts the Bernoulli likelihood (8) into a normal likelihood in terms of logit probabilities $\bm{\eta}$ conditional on Pólya-Gamma variables $\bm{\omega}$. Leveraging (5), $\pi(\bm{\eta}\mid\lambda)$ is also a multivariate normal in terms of $\bm{\eta}$. This leads to joint conditional $\pi(\lambda,\bm{\eta}\mid\bm{\omega},\mathbf{z})\propto\pi(\mathbf{z}\mid\bm{\eta},\bm{\omega})\pi(\bm{\eta}\mid\lambda)\pi(\lambda)$ that is multivariate normal in terms of $\bm{\eta}$. Thus, the computation of collapsed conditional $\pi(\lambda\mid\bm{\omega},\mathbf{z})=\int\pi(\lambda,\bm{\eta}\mid\bm{\omega},\mathbf{z})\mathrm{d}\bm{\eta}$ can integrate out $\bm{\eta}$ by exploiting normal-normal conjugacy, greatly improving mixing behavior. Furthermore, again by conjugacy, Step 3 of Algorithm 9 updates $\bm{\eta}$ from a multivariate normal; for example when correlation matrix $\mathbf{R}$ is full rank,
$\bm{\eta}\mid\bm{\omega},\lambda,\mathbf{z}\sim N_{n}\big{[}(\bm{\Omega}+\lambda^{-1}\mathbf{R}^{-1})^{-1}\{(\mathbf{z}-0.5\bm{1}_{n})+0.5(a-b)\mathbf{R}^{-1}\bm{1}_{n}\},(\bm{\Omega}+\lambda^{-1}\mathbf{R}^{-1})^{-1}\big{]}.$
Thus, the logistic-beta process offers significant computational benefits compared to multivariate beta constructions that rely on coordinate-wise updating (Trippa et al.,, 2011; DeYoreo and Kottas,, 2018) or Metropolis-Hastings in high-dimensional space (Arbel et al.,, 2016).
Sampling $\lambda$ from a collapsed conditional in Step 2 remains nontrivial and simple random walk Metropolis-Hastings would involve the calculation of Pólya density ${\pi_{\textsc{po}}}$ that is numerically unstable to evaluate. Based on Proposition 2.1, we propose a novel class of adaptive Pólya proposals that bypasses the evaluation of the Pólya density. Denote $\mathcal{L}(\lambda)=N_{n}\big{\{}\bm{\Omega}^{-1}(\mathbf{z}-0.5\bm{1}_{n});0.5\lambda(a-b)\bm{1}_{n},\lambda\mathbf{R}+\bm{\Omega}^{-1}\big{\}}$, which is proportional to $\int\pi(\mathbf{z}\mid\bm{\eta},\bm{\omega})\pi(\bm{\eta}\mid\lambda)\mathrm{d}\bm{\eta}$ in terms of $\lambda$. The adaptive Pólya proposal scheme selects the proposal distribution as $\mathrm{Polya}(a^{\prime},b^{\prime})$ with suitably chosen $(a^{\prime},b^{\prime})$ pair such that $a^{\prime}+b^{\prime}=a+b$. Then, the acceptance ratio for a candidate $\lambda^{\star}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$ becomes $\min(1,\alpha_{\mathrm{MH}})$ with
$\alpha_{\mathrm{MH}}=\frac{{\pi_{\textsc{po}}}(\lambda;a^{\prime},b^{\prime})}{{\pi_{\textsc{po}}}(\lambda^{\star};a^{\prime},b^{\prime})}\frac{{\pi_{\textsc{po}}}(\lambda^{\star};a,b)\mathcal{L}(\lambda^{\star})}{{\pi_{\textsc{po}}}(\lambda;a,b)\mathcal{L}(\lambda)}=\exp\{(\lambda-\lambda^{\star})(ab-a^{\prime}b^{\prime})/2\}\frac{\mathcal{L}(\lambda^{\star})}{\mathcal{L}(\lambda)}$
where the Pólya density terms are simplified by Proposition 2.1, bypassing direct evaluation of ${\pi_{\textsc{po}}}$. In terms of choosing the pair $(a^{\prime},b^{\prime})$ in an adaptive fashion, we use the moment matching method with a running average of $\lambda$. See Section S.2.2 of the Appendix for details.
When $n$ is large, sampling from $n$-dimensional multivariate normal in Step 3 becomes computationally expensive. If the correlation kernel is chosen based on $q\ll n$ normalized feature maps, $\mathbf{R}=\bm{\Phi}\bm{\Phi}^{\mathrm{\scriptscriptstyle T}}$ is a low-rank matrix, and the hierarchical representation (7) in Proposition 2.3 can be employed. Then, instead of sampling $n$-dimensional multivariate normal, Step 3 becomes sampling $\bm{\gamma}$ from $q$-dimensional multivariate normal with covariance matrix $\mathbf{V}_{\bm{\gamma}}=(\mathbf{I}_{q}+\lambda\bm{\Phi}^{\mathrm{\scriptscriptstyle T}}\bm{\Omega}\bm{\Phi})^{-1}$ and mean $\mathbf{V}_{\bm{\gamma}}\lambda^{1/2}\bm{\Phi}^{\mathrm{\scriptscriptstyle T}}\{\mathbf{z}-0.5\bm{1}_{n}-0.5\lambda(a-b)\bm{\omega}\}$ and setting $\bm{\eta}=0.5\lambda(a-b)\bm{1}_{n}+\lambda^{1/2}\bm{\Phi}\bm{\gamma}$, greatly reducing computational burden. In Section S.2.1 of the Appendix, we present further computational strategies based on scalable Gaussian process methods that preserve marginal variances.
The Algorithm 9 can be further extended to settings where beta parameters $(a,b)$ are assumed to be unknown and a hyperprior is placed on $(a,b)$. The same blocking strategy can be employed by jointly sampling $(a,b,\lambda)$ from the collapsed conditional using particle marginal Metropolis-Hastings (Andrieu et al.,, 2010). See Section S.2.3 of the Appendix for details.
@@ 4 Logistic-beta dependent Dirichlet processes
4.1 Definition and prior support
@@
The latent logistic-beta process model can be easily employed in more complex Bayesian nonparametric models. In this section, we introduce a logistic-beta dependent Dirichlet process for modeling dependent random probability measures. Random probability measures $\{G_{x}:x\in\mathscr{X}\}$ follow a logistic-beta dependent Dirichlet process (LB-DDP) with concentration $b$, correlation kernel $\mathcal{R}$, and independent atom processes $\{\theta_{h}(x):x\in\mathscr{X}\}$, $h\in\mathds{N}$ with $G^{0}_{x}$ the base measure, if
@@
$G_{x}(\cdot)=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x)\}\prod_{l<h}\left[1-\sigma\{\eta_{l}(x)\}\right]\Big{)}\delta_{\theta_{h}(x)}(\cdot),\,\,\,\,\eta_{h}(x)\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R}),$ (10)
@@
for $h\in\mathds{N}$. Since the stick-breaking ratios $V_{h}(x):=\sigma\{\eta_{h}(x)\}$ have $\mathrm{Beta}(1,b)$ marginals for any $x\in\mathscr{X}$ and are independent across $h=1,2\dots$, the random probability measure $G_{x}$ for any given $x\in\mathscr{X}$ marginally follows a Dirichlet process with concentration parameter $b$. The hyperparameter $b$ has a familiar interpretation, aiding in elicitation.
The prior introduced in equation (10) inherits many of the amenable properties common to dependent Dirichlet processes. For instance, under mild conditions, the topological support of the LB-DDP coincides with the space of collections of all probability measures whose support is included in the support of base measure indexed by $x$, as the following result shows.
Theorem 4.1.
Consider an LB-DDP with an atom process $\{\theta(x):x\in\mathscr{X}\}$ with support $\Theta$ that can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure. Then, the logistic-beta dependent Dirichlet process has full weak support.
The above statement follows from direct applications of the results described in Barrientos et al., (2012) and is a key requirement for posterior consistency. In the next subsection, we investigate further properties unique to the LB-DDP prior in equation (10).
4.2 Prior dependence properties
To gain a deeper understanding of the dependence structure induced by the LB-DDP, in this section we provide a full characterization of two key quantities: (i) the probability that two realizations $\vartheta\sim G_{x}$ and $\vartheta^{\prime}\sim G_{x^{\prime}}$ from two discrete random probability measures are equal, which we refer to as the tie probability, and (ii) the correlation between $G_{x}$ and $G_{x^{\prime}}$ evaluated at common set $B$. We begin by considering the single-atoms LB-DDP, where $\theta_{h}(x)\equiv\theta_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}G^{0}$ in (10) so that atoms do not vary across $x$. Such an example allows us to describe the dependence induced by the weights alone.
Theorem 4.2.
@@ 4.2 Prior dependence properties
Consider a single-atoms LB-DDP $\{G_{x}:x\in\mathscr{X}\}$ with concentration parameter $b$, correlation kernel $\mathcal{R}$, and non-atomic base measure $G^{0}$. Let $\mu(x,x^{\prime})=E[\sigma\{\eta(x)\}\sigma\{\eta(x^{\prime})\}]$ with $\eta(x)\sim{\textsc{lbp}}(1,b,\mathcal{R})$. Then, for any realizations $\vartheta\sim G_{x}$ and $\vartheta^{\prime}\sim G_{x^{\prime}}$, it holds that
$\operatorname*{pr}(\vartheta=\vartheta^{\prime}\mid G_{x},G_{x^{\prime}})={(1+b)}/\left\{2\mu(x,x^{\prime})^{-1}-(1+b)\right\}.$ (11)
Moreover, for any Borel set $B$, we have
$\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}={(1+b)^{2}}/\left\{2\mu(x,x^{\prime})^{-1}-(1+b)\right\}.$ (12)
The tie probability in equation (11) provides an intuitive summary of the dependence between $G_{x}$ and $G_{x^{\prime}}$ under general single-atoms DDP. Further multiplying such a quantity by $1+b$, one also obtains the correlation between random probability measures in equation (12). The quantities in (11) and (12) attain the maximum of $1/(1+b)$ and $1$, respectively, when $x=x^{\prime}$. Interestingly, their minimum is achieved when $\mu(x,x^{\prime})$ is minimized, though the value depends on $b$ due to the nontrivial correlation lower bound described in Corollary 2.1. We emphasize that the choice of $b$ affects the minimum possible tie probability and correlation, and within that range, the tie probability and correlation can be controlled by the correlation kernel $\mathcal{R}$. In Figure S.1.2 in the Appendix, we illustrate how the correlation in equation (12) changes as a function of the distance between $x$ and $x^{\prime}$ under the Matérn kernel.
@@
Figure 4: Analysis of the flexibility ranges induced by single-atoms DDPs (M1)–(M4) against the parameter $b$ (displayed in $\log_{2}$ scale). (Left) Ranges of $\operatorname*{corr}\{V_{h}(x),V_{h}(x^{\prime})\}$, the correlation between $\mathrm{Beta}(1,b)$ stick-breaking variables. (Center) Ranges of tie probability (11). (Right) Ranges of $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ (12).
To investigate the improved flexibility of the LB-DDP prior, we compare the greatest lower bounds of correlation between stick-breaking ratios, tie probability (11) and correlation between random probability measures (12) against other DDP constructions with different multivariate beta distributions. Specifically, we consider four different single-atoms dependent Dirichlet process models based on dependent beta stick-breaking ratios: (M1) LB-DDP; (M2) The process proposed by DeYoreo and Kottas, (2018); (M3) The process proposed by Nieto-Barajas et al., (2012); (M4) Copula-based dependent Dirichlet process. For M3, the minimum tie probability and correlation is achieved when $V_{h}(x)$ and $V_{h}(x^{\prime})$ are independent of each other. For M4, the minimum is achieved with counter-monotonic beta random variables; examples include Gaussian copula with perfect negative correlation. For completeness, we add brief descriptions of M2–M4 and their correlation lower bounds in Section S.1.3 of the Appendix.
Figure 4 illustrates how the range of the correlation between stick-breaking ratios, denoted as $\operatorname*{corr}\{V_{h}(x),V_{h}(x^{\prime})\}$, affects the ranges of tie probabilities and correlation between the random measure. We remark that, due to single-atoms assumption, $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ does not depend on $B$ and is always positive regardless of $\operatorname*{corr}\{V_{h}(x),V_{h}(x^{\prime})\}$; see Ascolani et al., (2024). The copula-based formulation M4 achieves the highest level of flexibility, which is an expected result due to Sklar’s theorem (Joe,, 2014). However, inference with M4 often relies on Metropolis-Hastings in potentially high-dimensional space (Arbel et al.,, 2016), and the gap between M1 and M4 in Figure 4 corresponds to the cost of expressiveness that single-atoms LB-DDP pays to achieve computational tractability. The behavior of the minimum tie probability and correlation of M1 varies distinctly under different values of $b$ compared to others. For small to moderate $b$, M1 can capture a much wider range of dependence compared to M2 and M3 by allowing negatively correlated stick-breaking ratios. When $b$ is large, the range of M1 becomes limited due to the nontrivial lower bound of correlation described in Corollary 2.1. This suggests that when $b>8$, the single-atoms assumption needs to be relaxed if one wants to ensure maximum flexibility relative to the independent weight model (infimum under M3). We formalize this in the following statement, providing a more general result of Theorem 4.2.
Theorem 4.3.
Consider the same settings as Theorem 4.2, but the single-atoms assumption is relaxed with i.i.d. (across $h$) atom processes $\{\theta_{h}(x):x\in\mathscr{X}\}$ with $G^{0}$ marginals for any $x$. Let $G^{0}_{x,x^{\prime}}(B)=\operatorname*{pr}(\theta_{h}(x)\in B,\theta_{h}(x^{\prime})\in B)$ which does not depend on $h$. Then,
$\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}=\rho_{0}(B)\times{(1+b)^{2}}/\left\{2\mu(x,x^{\prime})^{-1}-(1+b)\right\},$ (13)
where $\rho_{0}(B)=\{G^{0}_{x,x^{\prime}}(B)-G^{0}(B)^{2}\}/\{G^{0}(B)-G^{0}(B)^{2}\}$. Moreover, $\rho_{0}(B)\in[-1,1]$.
For example, when the atom process is positively correlated but not identical across $x$, we have $G^{0}(B)^{2}<G^{0}_{x,x^{\prime}}(B)<G^{0}(B)$ so that $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}$ is multiplied by a fractional factor $\rho_{0}(B)$ compared to the single-atoms DDP case, alleviating the restricted range of correlation when $b$ is large.
In addition to the study of correlation structure, ensuring the continuity of the process as a function of covariates is important in the development of dependent Bayesian nonparametric processes (MacEachern,, 2000). In Theorem 4.4, we show that the continuity of correlation kernel $\mathcal{R}$ at the diagonal leads to the continuity of the single-atoms LB-DDP as a function of a covariate in terms of correlation between random probability measures.
Theorem 4.4.
Consider a single-atoms LB-DDP same as the setting of Theorem 4.2, and let $\mathcal{X}$ be a compact subset of $\mathbb{R}^{d}$. If the correlation kernel $\mathcal{R}(x,x^{\prime})$ is continuous at every diagonal point $x=x^{\prime}$, then $\operatorname*{corr}\{G_{x}(B),G_{x^{\prime}}(B)\}\to 1$ as $x\to x^{\prime}$.